OpenStudy (anonymous):

Please Help!!? Find if G o F if f(x) = x + 7 and g(x) = 3x – 2. A. 3x + 5 B. 3x - 5 C. 3x + 19 D. -3x - 19

OpenStudy (anonymous):

What do you mean by: G o F ?

OpenStudy (anonymous):

g(f(x))=g(x+7)=3(x+7)-2 =3x+21-2 =3x+19 Choice c is the correct answer